Timeline of Mathematicians
1700 B.C.E.
Ahmes (c. 1650 B.C.E.)
700 B.C.E.
600 B.C.E.
Thales of Miletus (c. 630-c 550)
Pythagoras of Samos (c. 570-c. 490)
500 B.C.E.
Socrates (469-399)
400 B.C.E.
Plato (427-347)
Eudoxus of Cnidos (c. 400-c. 347)
Aristotle (384-322)
300 B.C.E.
Euclid (fl. c. 295)
Archimedes of Syracuse (287-212)
Eratosthenes of Cyrene (c. 276-c. 195)
100 B.C.E.
Zeno of Sidon (c. 79 BCE)
100 C.E.
Ptolemy (Claudius Ptolemaeus) (c. 100-c. 170)
200 C.E.
Diophantus of Alexandria (c. 250?)
300 C.E.

Pappus of Alexandria (fl. c. 300-c. 350)
Hypatia of Alexandria (c. 370-415)
900
Abu l-Quasim Maslama ibn Ahmad al-Faradi al-Majriti (fl. 980-1000)
1000
`Umar al-Khayyami (Omar Khayyam) (c. 1048-c. 1131)
1100
Leonardo Fibonacci of Pisa (C. 1170-post 1240)
1300
William of Ockham (c. 1285-c. 1349)
1400
Piero della Francesca (c. 1410-1492)
Leonardo da Vinci (1452-1519)
Scipione del Ferro (1465-1526)
Nicolas Copernicus (1473-1543)
1500
Girolamo Cardano (1501-1576)
Robert Recorde (1510-1558)
Gerardus Mercator (Kremer) (1512-1594)
Ludovico Ferrari (1522-1565)
1550
Franois Vite (Vieta) (1540-1603)
Ludolph van Ceulen (1540-1610)
John Napier (1550-1617)
Francis Bacon (1561-1626)
Henry Briggs (1561-1631)
Galileo Galilei (1564-1642)
1600
Ren du Perron Descartes (1596-1650)

Pierre de Fermat (1601-1665)
John Wallis (1616-1703)
Nicolas Mercator (Kaufman) (1620-1687)
Blaise Pascal (1623-1662)
1650
Isaac Barrow (1630-1677)
James Gregory (1638-1675)
Isaac Newton (1642-1727)
Gottfried Wilhelm Leibniz (1646-1716)
1675
Jacques Bernoulli (James, Jakob) (1654-1705)
Jean Bernoulli (John, Johann) (1667-1748)
1700
Nicholas Bernoulli (1687-1759) (nephew of Jean)
Christian Goldbach (1690-1764)
1725
Daniel Bernoulli (1700-1782)
Leonhard Euler (1707-1783)
1750
Johann Heinrich Lambert (1728-1777)
Edward Waring (1734-1798)
Joseph Louis Lagrange (1736-1813)
Jean-Baptiste Joseph Delambre (1749-1822)
Pierre Simon de Laplace (1749-1827)
1775
Adrien-Marie Legendre (1752-1833)
Jean Baptiste Joseph Fourier (1768-1830)
1800
Carl Friedrich Gauss (1777-1855)
August Leopold Crelle (1780-1855)

William Hamilton (1788-1856)
Augustin-Louis Cauchy (1789-1857)
Augusus Ferdinand Mbius (1790-1868)
George Green (1793-1841)
1820
Niels Henrik Abel (1802-1829)
Carl Gustav Jacob Jacobi (1804-1851)
1830
William Rowan Hamilton (1805-1865)
Benjamin Peirce (1809-1880)
Joseph Liouville (1809-1882)
variste Galois (1811-1832)
Auguste Bravais (1811-1863)
James Joseph Sylvester (1814-1897)
1840
Ada Lovelace (1815-1852)
Karl Weierstrass (1815-1897)
Pafnuty Lvovich Chebyshev (1821-1894)
Charles Hermite (1822-1901)
Zacharias Dase (1824-1861)
1850
Georg Friedrich Bernhard Riemann (1826-1866)
1860
Franois Marius Sophus Lie (1842-1899)
Victor Schlegel (1843-1905)
1870
Georg Cantor (1845-1918)
Christian Felix Klein (1849-1925)
Carl Louis Ferdinand Lindemann (1852-1939)
Jules Henri Poincar (1854-1912)

Thomas Jan Stieltjes (1856-1894)
Andrei Andreyevich Markov (1856-1922)
Max Karl Ernst Ludwig Planck (1858-1947)
John Charles Fields (1863-1932)
1890
Jacques Hadamard (1865-1963)
Charles-Jean de la Valle Poussin (1866-1962)
Ernst Friedrich Ferdinand Zermelo (1871-1951)
Bertrand Russell (1872-1970)
Leonard Eugene Dickson (1874-1954)
1900
Albert Einstein (1879-1955)
Otto Szsz (1884-1952)
1910
Srinivasa Aaiyangar Ramanujan (1887-1920)
George Plya (1887-1985)
1920